OpenStudy (anonymous):

Can anyone get the derivative of this ? y = 1/3 Log e^(3x+4) + Log e^(6x+8)

zepdrix (zepdrix):

Are those logs of base 10? Or is it like this? \(\Large\rm \log_e (3x+4)\) The notation is kind of weird... so I can't tell..

OpenStudy (anonymous):

It a relief you'd answer me .

zepdrix (zepdrix):

Need clarification on the format. This: \(\Large\rm y=\frac{1}{3}\log e^{3x+4}+\log e^{6x+8}\) Or this: \(\Large\rm y=\frac{1}{3}\log_e(3x+4)+\log_e(6x+8)\) Which is the correct format? The "^" is confusing.

OpenStudy (anonymous):

If its ok Can you solve them Both . My teacher did not clarify it thats why when I simplified it to In(3x+4) +In(6x+8) they said it is wrong

zepdrix (zepdrix):

Let's start with the second one, that one makes more sense.\[\Large\rm y=\frac{1}{3}\log_e(3x+4)+\log_e(6x+8)\]\[\Large\rm y=\frac{1}{3}\ln(3x+4)+\ln(6x+8)\]Recall the derivative of ln(x) is 1/x. Same idea here, derivative of ln(stuff) is 1/(stuff) then we have to remember to apply the chain rule as well.

OpenStudy (anonymous):

Yeah That What I did . Then I got The answer y ' =1/ (3x+4) + 6/ (6x+8)

zepdrix (zepdrix):

\[\Large\rm y'=\frac{1}{3}\frac{1}{3x+4}\color{royalblue}{(3x+4)'}+\frac{1}{6x+8}\color{royalblue}{(6x+8)'}\] \[\Large\rm y'=\frac{1}{3}\frac{1}{3x+4}\color{royalblue}{(3)}+\frac{1}{6x+8}\color{royalblue}{(6)}\]And that was incorrect? Ok let's try the other form then.

zepdrix (zepdrix):

We'll assume these are logs of base 10,\[\Large\rm y=\frac{1}{3}\log \left(e^{3x+4}\right)+\log \left(e^{6x+8}\right)\] A rule of logs:\[\large\rm \log\left(a^{\color{orangered}{b}}\right)=\color{orangered}{b}\log(a)\]Applying this rule gives us:\[\Large\rm y=\frac{1}{3}(3x+4)\log (e)+(6x+8)\log (e)\]

OpenStudy (anonymous):

Then Find its derivative now ?

zepdrix (zepdrix):

Yes. Log base 10 of e is just a CONSTANT. Don't let it confuse you! We don't have to do any fancy product rule or anything like that.

zepdrix (zepdrix):

\[\Large\rm y=\frac{\log(e)}{3}(3x+4)+\log(e)(6x+8)\]It's in the form:\[\Large\rm y=A(3x+4)+B(6x+8)\]Where A and B are constants. So just differentiate what is inside of the brackets.

OpenStudy (anonymous):

The after differentiating the thing inside the parenthesis . whats next that I need to do ?

zepdrix (zepdrix):

\[\Large\rm y'=\frac{\log(e)}{3}(3)+(6)\log(e)\]Mmm nothing really :\ Just simplify the 3's. I doubt this is the correct form of the question though. It's really confusing, not knowing what they're actually asking.

OpenStudy (anonymous):

Is the First solution which is The answer 1 / (3x+4) + 6 / (6x+8) . Is correct base on the BASE e or not ?

zepdrix (zepdrix):

Yes, our first solution corresponded to an equation containing logs of base e.

OpenStudy (anonymous):

\[y' =\frac{ 1 }{ 3x+4 }+ \frac{ 6 }{ 6x+8}\]

OpenStudy (anonymous):

Is that Correct ?

zepdrix (zepdrix):

Yes

OpenStudy (anonymous):

So , Mr. Zepdrix What is the Answer on the second situation ? Im sorry im giving you a hard time on those things that makes my mind confuse . But Last week My prof . Give a unsure Answer . Which is 7loge I dont know if its correct or not . My prof also did not know If its correct.

zepdrix (zepdrix):

For the second form, we assumed the logs were of base 10 and applied a rule of logs before differentiating. Then after taking our derivative, and simplifying, we end up with:\[\Large\rm y'=\log_{10}(e)+6\log_{10}(e)\]

OpenStudy (anonymous):

Its also 7log e Right with Base of 10 ?

zepdrix (zepdrix):

Oh yes yes. Like-terms, we can combine! :) Correct.

OpenStudy (anonymous):

Thank you so much Btw im taking Diff Cal. now in school Thats Why I ask you about this sort of Questions.

zepdrix (zepdrix):

Oh cool c: Calculus is so much fun.

OpenStudy (anonymous):

Do you have any website that will help me improve in Math ?

zepdrix (zepdrix):

Khan Academy is a great resource. https://www.khanacademy.org I used this website a ton before taking Calculus. https://www.math.ucdavis.edu/~kouba/ProblemsList.html It really helped me to understand things and get ready for the subject. MIT has their OpenCourseWare on Youtube, it's really excellent. https://www.youtube.com/watch?v=7K1sB05pE0A&list=PL590CCC2BC5AF3BC1 https://www.youtube.com/watch?v=PxCxlsl_YwY&list=PL4C4C8A7D06566F38 https://www.youtube.com/watch?v=XDhJ8lVGbl8&list=PLEC88901EBADDD980

OpenStudy (anonymous):

Wow Great!! Thank you so much